GAS-SURFACE DYNAMICS OF VIBRATIONALLY AND ROTATIONALLY EXCITED REAGENTS: METHANE (ν_{3}, J=0-3 AND 3ν_{4}, J=2) DISSOCIATION ON NICKEL SURFACES

\underline{A. L. UTZ}, R. R. SMITH, AND D. R. KILLELEA, Department of Chemistry, Tufts University, Medford, MA, 02155; L. B. F. JUURLINK, Department of Chemistry, Leiden University.

A high resolution infrared laser prepares a significant population of vibrationally excited methane molecules in a supersonic molecular beam. The prepared molecules impinge on a clean Ni(100) or Ni(111) surface housed in an ultrahigh vacuum chamber. Some of the methane dissociates and chemically binds to the surface upon impact. We quantify reaction probability as a function of incident translational energy, Ni surface temperature, and methane’s vibrational and rotational state. The state-resolved reaction probabilities reveal how energy deposited into different energetic coordinates activates methane’s dissociative chemisorption. We find that the more energetic 3ν_{4} bending state is less reactive than ν_{3} antisymmetric C-H stretch state on both Ni(100) and Ni(111). This result highlights the presence of vibrational mode specificity in a gas-surface reaction on a metal surface. We also find that vibrational energy in ν_{3} is more effective than translational energy in its ability to promote reaction on Ni(111). Both of these results point to the importance of non-statistical energy flow in this reaction.